OpenStudy (anonymous):

Simplify each term, if necessary, and then use the distributive property to combine terms, if possible.

OpenStudy (anonymous):

\[\sqrt{108}+2\sqrt{12}-\sqrt{3}\]

OpenStudy (unklerhaukus):

factorise the terms under square roots into products of prime numbers

OpenStudy (unklerhaukus):

what are the prime factors of 108?

OpenStudy (anonymous):

\[\sqrt{108}=6\sqrt{3}\]

OpenStudy (anonymous):

\[2\sqrt{12}=2(2\sqrt{3})=4\sqrt{3}?\]

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

I did it !

OpenStudy (unklerhaukus):

that is good, \[\sqrt{108}+2\sqrt{12}-\sqrt{3}\] \[=6\sqrt{3}+4\sqrt{3}-\sqrt{3}\] now you can the distributive property \[ax+bx+cx=(a+b+c)x\]